# Interaction induced reentrance of Bose glass and quench dynamics of Bose gases in twisted bilayer and quasicrystal optical lattices

### v1

@@ Abstract
We investigate the ground state and dynamical properties of ultracold gases in optical lattices with an aperiodic external potential—a scenario motivated by recent experiments on twisted bilayer optical lattices and optical quasicrystals. Our study reveals that the interplay between on-site repulsive interactions and a quasiperiodic potential gives rise to rich physics. At low filling factors, increasing the interaction strength induces a delocalization effect that transforms a Bose glass (BG) phase, characterized by disconnected superfluid (SF) regions, into a robust SF phase with a percolated network of SF clusters. This transition is quantitatively characterized by monitoring the percolation probability. At higher filling factors, we uncover a striking reentrant behavior: as the on-site interaction increases, the system initially transitions from BG to SF, but a further increase reverses this trend, returning the system to the BG phase. This reentrance is ascribed to an interaction-driven rearrangement of particles, where a once percolated SF network fragments into isolated SF islands as repulsive interactions dominate. Furthermore, our analysis of quench dynamics demonstrates distinct transient behaviors. Intra-phase quenches yield minimal variations in both the percolation probability and the inverse participation ratio (IPR) of the particle density distribution. In contrast, inter-phase quenches produce pronounced effects; for instance, a quench from the SF to BG phase is marked by an abrupt loss of global SF connectivity, while a BG-to-SF quench features oscillatory changes in the percolation probability and a gradual decrease in the IPR, eventually stabilizing the SF phase. Our findings unveil the complex interplay between disorder and interaction in ultracold Bose gases, offering valuable insights that are highly pertinent to current experimental efforts employing state-of-the-art twisted bilayer and quasicrystalline optical lattice platforms.
@@ I Introduction
Recent years, twisted bilayer materials opens a brand new era of twistronics, which has become a fertile soil for studying and engineering strongly correlation phenomena (Cao et al., 2018a; Li et al., 2021a; Cao et al., 2018b; Gong et al., 2017; Huang et al., 2017; Chen et al., 2020; Li et al., 2021b; Regan et al., 2020; Serlin et al., 2020; Li et al., 2021c; Zhou et al., 2021). Most researches focus on the moiré patterns, the commensurate patterns of multiple twisted crystal layers, where the whole system remains certain translational symmetry but the unit cell is much larger than the original one, leading to the flattening of bands and thus, enhancing the interaction effects. Experimentalists have observed in laboratory many emergent strongly correlation effects, such as correlated insulators (Cao et al., 2018a; Li et al., 2021a), superconductivity (Cao et al., 2018b), magnetism (Gong et al., 2017; Huang et al., 2017; Chen et al., 2020), generalized Wigner crystals (Li et al., 2021b; Regan et al., 2020), integer and fractional quantum anomalous Hall insulators (Serlin et al., 2020; Li et al., 2021c; Zhou et al., 2021; Cai et al., 2023; Park et al., 2023; Zeng et al., 2023; Xu et al., 2023; Lu et al., 2024), etc.
@@
On the other hand, the quasicrystal patterns can emerge when the crystal layers are twisted in an incommensurate way that completely breaks the translational symmetry but preserves the long-range order (Shechtman et al., 1984), thereby mediating between the order and disorder. The quasicrystals have been experimentally realized and highly tuned by twisting three layers of graphenes (Uri et al., 2023) or bilayers of tungsten diselenides (Li et al., 2024), but has been less focused before (Park et al., 2019; Huang and Liu, 2019) than the commensurate case. It is well known that a quasicrystal would lead to the Anderson localization (Anderson, 1958), a phenomenon that lacks the diffusiveness of non-interacting particles when subjected to a random or quasiperiodic potential landscape. In one-dimensional (1D) and two-dimensional (2D) cases, all single-particle eigenstates can be localized by a finite strength of quasicrystal potential (Aubry and André, 1980; Roati et al., 2008; Lahini et al., 2009), different from the random disorder where the localization occurs for any infinitesimal strength (Abrahams et al., 1979). Therefore, quasicrystals are ideal platforms for studying localization physics.
@@
Now that the moiré pattern can induce many exotic strongly correlated phenomena, one may wonder how the localization feature in twisted bilayer quasicrystals competes with the interaction. In fermionic systems, most researches focus on the thermal properties of many body localization, especially in 1D case (Iyer et al., 2013; Mondaini and Rigol, 2015; Xu et al., 2019; Vu et al., 2022; Wang et al., 2021). And the incommensurability is also found to radically change the ground state of interacting fermions, compared with the commensurate case (Vu and Das Sarma, 2021; Gonçalves et al., 2024a; Kraus et al., 2014; Cookmeyer et al., 2020; Oliveira et al., 2024; Gonçalves et al., 2024b). In bosonic systems, the interplay of the randomness/quasiperiodicity and the interaction would lead to an exotic many-body ground state—Bose glass (BG) (Giamarchi and Schulz, 1988; Fisher et al., 1989), which is insulating like an Anderson insulator, but compressible like a superfluid. Disordered interacting bosons in different dimensions have been studied in various experimental platforms, such as helium films on porous substrates (Crowell et al., 1997), disordered superconducting films (Sacépé et al., 2011), doped quantum magnets (Yu et al., 2012), and especially in optical lattices (Fallani et al., 2007; Deissler et al., 2010; Pasienski et al., 2010; Gadway et al., 2011; D’Errico et al., 2014; Meldgin et al., 2016). There are also plenty of theoretical studies for interacting bosons in disordered or quasiperiodic potentials (Rapsch et al., 1999; Roux et al., 2008; Bissbort and Hofstetter, 2009; Söyler et al., 2011; Niederle and Rieger, 2015; Zhang et al., 2015; Gerster et al., 2016; Yao et al., 2020; Johnstone et al., 2021), and even bosons in interaction-induced disordered or quasiperiodic potentials (Zeng et al., 2024). However, it is still lacking for the study of interacting bosons in quasicrystals of twisted bilayer optical lattices. Recently, Yu et al. experimentally created a 2D octagonal quasicrystalline lattice using four light beams in a plane and one beam in the perpendicular direction, realizing a BG with ultracold atoms in the weakly interacting regime (Yu et al., 2024; Gautier et al., 2021).
@@
In this work, we study the ground state and dynamical properties of a class of ultracold bosonic systems realized in optical lattices with an aperiodic external potential—a setting directly inspired by recent experiments on twisted bilayer optical lattices (Meng et al., 2023) and optical quasicrystals (Yu et al., 2024). We investigate how the interplay between on-site repulsive interactions and a quasiperiodic potential drives the system between a homogeneous superfluid (SF) phase and a BG phase. At low filling factors, our findings reveal that an increase in the on-site interaction strength induces a delocalization effect, transforming a BG phase—with disconnected SF regions lacking global connectivity—into a robust SF phase characterized by a percolated network of SF clusters. This phase transition is quantitatively characterized by monitoring the percolation probability (see Fig. 1). At higher filling factors, our results uncover a striking reentrant behavior (see Fig. 2): the system initially undergoes a transition from a BG to an SF phase as the on-site interaction strength increases, but a further increase in the interaction leads to a reversal back to the BG phase. This reentrance is attributed to the interaction-driven rearrangement of particles, wherein a once percolated SF network eventually fragments into isolated SF islands as repulsive interactions become dominant. Moreover, our examination of the system’s quench dynamics reveals distinct transient behaviors (see Fig. 3): while intra-phase quenches produce minimal changes in both the percolation probability and inverse participation ratio (IPR) of the particle density distribution, inter-phase quenches lead to pronounced dynamical effects. Specifically, a quench from the SF to BG phase results in an abrupt loss of global SF connectivity, whereas a BG-to-SF quench is marked by oscillatory changes in the percolation probability and a gradual decrease in the IPR, ultimately stabilizing the SF phase. These findings unveil the rich physics of ultracold Bose gases in optical lattices with an aperiodic external potential, offering valuable insights that are highly relevant for state-of-the-art experiments employing twisted bilayer or quasicrystalline optical lattice platforms.
@@ II System and Model
Inspired by the recent experimental progresses in realizing twisted bilayer optical lattice (Meng et al., 2023) and optical quasicrystal (Yu et al., 2024) for ultracold atoms, we consider a single component ultracold bosonic system, the Hamiltonian of which takes the form of a “single-layer” Bose-Hubbard model with an additional external on-site potential, i.e.,
@@
[…]
@@
where $M_{r}$ is the strength of the external potential, $i_{x}$ ($i_{y}$) is the index of the $i\textrm{th}$ site of the square lattice along the $x$ ($y$) direction, and we choose $\theta=45\text{\textdegree}$. This corresponds to the case in the bilayer systems with the twist-angle being $45\text{\textdegree}$ (Meng et al., 2023), or the one in the optical quasicrystal system with the optical lattice formed by superimposing four independent one-dimensional lattices in the $x$-$y$ plane at $45\text{\textdegree}$ angle (Yu et al., 2024).
The first two terms of the Hamiltonian (1) assume the form of the conventional Bose-Hubbard model (Fisher et al., 1989), they favor two homogeneous phases, namely, the homogeneous SF phase (for large $J/U$) and Mott insulator phase (for small $J/U$ at integer filling factors) that respect the discrete translation symmetry of the underlying lattice. The third term of $\hat{H}$ is an aperiodic or a quasiperiodic external potential that explicitly breaks the lattice translation symmetry. Recent investigations (Johnstone et al., 2021) showed that this type of terms can give rise to the BG phase that usually arises in disordered bosonic systems (Fisher et al., 1989), due to their aperiodicity. In fact, the interplay between on-site interaction and the quasiperiodic external potential can give rise to much richer physics associated to the BG phase. Particularly, at fixed particle densities which is highly relevant for corresponding experiments, this interplay can give rise to interaction induced reentrant transitions for BG phases and rich quench dynamics as we shall show in the following.
@@ III Results
To study the ground state of the system described by Hamiltonian (1), we used the bosonic Gutzwiller variational method (Krauth et al., 1992; Jaksch et al., 1998; Lanatà et al., 2012). The variational ground state is assumed to take the site-factorized form $\left|\textrm{GW}\right\rangle=\left|\phi_{1}\right\rangle\otimes\cdots\otimes\left|\phi_{i}\right\rangle\otimes\cdots\otimes\left|\phi_{N_{\textrm{lat}}}\right\rangle$, where each $\left|\phi_{i}\right\rangle$ represents the state at the $i$-th lattice site, and $N_{\textrm{lat}}$ denotes the total number of sites in the lattice. The individual single-site states $\left|\phi_{i}\right\rangle$ are expanded as a superposition of particle-number eigenstates as $\left|\phi_{i}\right\rangle=\sum_{n=0}^{n_{\textrm{max}}}C_{i.n}\left|n\right\rangle_{i}$, where $n_{\textrm{max}}$ is the maximum number of particles per site and $C_{i.n}$ are variational parameters that are to be determined by minimizing the total energy of the system. In this work, we focus on the ground state properties of the system at different filling factors $\rho\equiv N/N_{\textrm{lat}}$, with $N$ being the total number of particles in the system. We examine the system at different filling factors to explore how particle density influences the ground state properties. For the numerical calculations, we set square lattice size $L=51$, the local maximum truncation number $n_{\textrm{max}}=10$, and we employ open boundary conditions for the numerical results if not specified in the text.
Figure 1: Phase diagram and representative real-space distributions. (a) Phase diagram for a system of $100$ particles on a $51\times 51$ square lattice, where the black solid line is the phase boundary between BG (upper) and SF (lower) phases. Inset: Dependence of the percolation probability $\mathcal{P}$ (blue curve) and the inverse participation ratio (IPR) (red curve) on $U/J$ for points along the dotted line indicated in the phase diagram. The inserted SF order parameter distributions correspond to the nearby red points marked in the figure. (b) Real-space distributions of the particle density $\left\langle n_{i}\right\rangle$, SF order parameter $\left\langle b_{i}\right\rangle$ and the discrete field $S_{i}$ near the phase boundary ($M_{r}/J=5.9$). See text for more details.
@@ III.1 Transitions between superfluids and Bose glasses
At relatively low filling factors, the typical properties of the system are summarized in Fig. 1, which shows the phase diagram of the system at $\rho\approx 0.04$, along with the particle density distribution $\left\langle\hat{n}_{i}\right\rangle$, the SF order parameter distribution $\langle\hat{b}_{i}\rangle$, and the discrete field distribution $S_{i}$. The latter is used to distinguish between SF and BG phases and will be discussed in detail later. As shown in the lower-left inset of Fig. 1(a), in the low aperiodic potential regime, as expected from the homogeneous Bose-Hubbard model, the system is in the SF phase, with a nonzero SF order parameter $\langle\hat{b}_{i}\rangle$ that is distributed relatively evenly across the entire lattice. In contrast, in the high aperiodic potential regime, as shown in the upper inset of Fig. 1(a), the system maintains a non-zero SF order parameter $\langle\hat{b}_{i}\rangle$, but its distribution exhibits a qualitatively different structure. In particular, although the SF order parameter is nonzero in various regions, these regions are “disconnected” from each other, indicating a transition to a new phase: the BG phase, which is absent in conventional homogeneous systems of lattice bosons.
@@
To quantitatively distinguish the SF phase from the BG phase, we employ the percolation method (Sheshadri et al., 1995; Buonsante et al., 2009; Kemburi and Scarola, 2012; Niederle and Rieger, 2013; Johnstone et al., 2021) , a widely used technique for studying phase transitions in disordered or quasi-disordered systems. The key quantity used to differentiate the SF and BG phases is the percolation probability $\mathcal{P}$, with $\mathcal{P}=0$ indicating no connected SF region spanning the entire system (signifying the BG phase), and $\mathcal{P}>0$ indicating the presence of a percolated SF region (signifying the SF phase spanning the entire system). In this work, we adopt the method from (Niederle and Rieger, 2013; Johnstone et al., 2021), where $\mathcal{P}$ is calculated using the corresponding discrete field $S_{i}$ (Niederle and Rieger, 2013), which is constructed from both the particle density distribution $\left\langle\hat{n}_{i}\right\rangle$ and the SF order parameter distribution $\langle\hat{b}_{i}\rangle$. This field encapsulates the percolation properties of the SF regions in the system (see Appendix A for technical details on extracting $\mathcal{P}$ from $S_{i}$). For example, in the last row of Fig. 1(b), red-marked sites represent various connected SF regions that do not span the entire system, while blue-marked sites correspond to connected SF regions that do span the entire system, forming a percolated SF region.
@@
From Fig. 1(a), we observe that tuning the on-site interaction strength $U$ can drive the transition between the SF and BG phases in the high aperiodic potential regime. As indicated by the horizontal arrow in Fig. 1(a) and the corresponding distributions for $\left\langle\hat{n}_{i}\right\rangle$, $\langle\hat{b}_{i}\rangle$, and $S_{i}$ in Fig. 1(b) at a fixed aperiodic potential strength $M_{r}$, the system transitions from the BG phase with $\mathcal{P}=0$ at relatively small $U$ to the SF phase with a finite percolation probability $\mathcal{P}$ at larger $U$ [see the lower-right inset of Fig. 1(a)]. This transition is driven by the “delocalization” effect of the repulsive on-site interaction. As seen in the lower-right inset of Fig. 1(a), the IPR for the normalized density distribution $\widetilde{\left\langle\hat{n}_{i}\right\rangle}\equiv\left\langle\hat{n}_{i}\right\rangle/\sum_{i=1}^{N_{\textrm{lat}}}\left\langle\hat{n}_{i}\right\rangle$ directly characterizes the extent of localization in the system, i.e., $\textrm{IPR}\equiv\sum_{i=1}^{N_{\textrm{lat}}}\widetilde{\left\langle\hat{n}_{i}\right\rangle}^{2}$. The IPR decreases as $U$ increases, reflecting the delocalization of the particles. At relatively small $U$, the SF regions are localized [see the left most plot for $S_{i}$ in Fig. 1(b)]. As $U$ increases, the SF regions expand, and at large enough $U$, they can merge, forming a percolated SF region that drives the system into the SF phase.
@@
Figure 2: Phase diagram and representative real-space distributions. (a) Phase diagram for a system of $196$ particles on a $51\times 51$ square lattice, where the black solid line is the phase boundary between BG (upper) and SF (lower) phases. Inset: Dependence of the percolation probability $\mathcal{P}$ (blue curve) and IPR (red curve) on $U/J$ for points along the dotted line indicated in the phase diagram. (b) Real-space distributions of the particle density $\left\langle n_{i}\right\rangle$, SF order parameter $\left\langle b_{i}\right\rangle$ and the discrete field $S_{i}$ near the phase boundary ($M_{r}/J=8$). See text for more details.

### v3

# Interaction-induced reentrance of Bose glass and quench dynamics of Bose gases in twisted bilayer and quasicrystal optical lattices

@@ Abstract
We investigate the ground state and dynamical properties of ultracold gases in optical lattices with a quasicrystal structure—a scenario inspired by recent experiments on twisted bilayer optical lattices and optical quasicrystals. Our study reveals that the interplay between on-site repulsive interactions and a quasiperiodic potential gives rise to rich physics. At low filling factors, increasing the interaction strength induces a delocalization effect that transforms a Bose-glass (BG) phase, characterized by disconnected superfluid (SF) regions, into a robust SF phase with a percolated network of SF clusters. This transition is quantitatively characterized by monitoring the percolation probability. At higher filling factors, we uncover a striking reentrant behavior: As the on-site interaction increases, the system initially transitions from BG to SF, but a further increase reverses this trend, returning the system to the BG phase. This reentrance is ascribed to an interaction-driven rearrangement of particles, where a once percolated SF network fragments into isolated SF islands as repulsive interactions dominate. Furthermore, our analysis of quench dynamics demonstrates distinct transient behaviors. Intra-phase quenches yield minimal variations in both the percolation probability and the inverse participation ratio (IPR) of the particle density distribution. In contrast, interphase quenches produce pronounced effects; for instance, a quench from the SF to BG phase is marked by an abrupt loss of global SF connectivity, while a BG-to-SF quench features oscillatory changes in the percolation probability and a gradual decrease in the IPR, eventually stabilizing the SF phase. Our findings unveil the complex interplay between quasiperiodic optical lattice potential and interaction in ultracold Bose gases, offering valuable insights that are highly pertinent to current experimental efforts employing state-of-the-art twisted bilayer and quasicrystalline optical lattice platforms.
@@ I Introduction
Recently, twisted bilayer materials have opened a brand-new era of twistronics, which has become fertile ground for studying and engineering strong-correlation phenomena [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11]. Most research has focused on moiré patterns, the commensurate patterns of multiple twisted crystal layers, in which the whole system retains certain translational symmetry but the unit cell is much larger than the original one, leading to the flattening of bands and thus enhancing interaction effects. Experimentalists have observed in the laboratory many emergent strong-correlation effects, such as correlated insulators [1, 2], superconductivity [3], magnetism [4, 5, 6], generalized Wigner crystals [7, 8], and integer and fractional quantum anomalous Hall insulators [9, 10, 11, 12, 13, 14, 15, 16].
@@
On the other hand, quasicrystal patterns can emerge when the crystal layers are twisted in an incommensurate way that completely breaks the translational symmetry but preserves the long-range order [17], thereby mediating between order and disorder. Quasicrystals have been experimentally realized and highly tuned by twisting three layers of graphene [18] or bilayers of tungsten diselenides [19] but have received less attention before [20, 21] than the commensurate case. It is well-known that a quasicrystal will lead to Anderson localization [22], a phenomenon that lacks the diffusiveness of noninteracting particles when it is subjected to a random or quasiperiodic potential landscape. In the one-dimensional (1D) and two-dimensional (2D) cases, all single-particle eigenstates can be localized by a finite strength of the quasicrystal potential [23, 24, 25], different from the random disorder where the localization occurs for any infinitesimal strength [26]. Therefore, quasicrystals are ideal platforms for studying localization physics.
@@
Now that the moiré pattern can induce many exotic strongly correlated phenomena, one may wonder how the localization feature in twisted bilayer quasicrystals competes with the interaction. In fermionic systems, most research has focused on the thermal properties of many-body localization, especially in the 1D case [27, 28, 29, 30, 31]. The incommensurability is also found to radically change the ground state of interacting fermions, compared with the commensurate case [32, 33, 34, 35, 36, 37]. In bosonic systems, the interplay of the randomness or quasiperiodicity and the interaction will lead to an exotic many-body ground state—Bose glass (BG) [38, 39], which is insulating like an Anderson insulator but compressible like a superfluid. Disordered interacting bosons in different dimensions have been studied using various experimental platforms, such as helium films on porous substrates [40], disordered superconducting films [41], doped quantum magnets [42], and, especially, optical lattices [43, 44, 45, 46, 47, 48]. There are also plenty of theoretical studies on interacting bosons in disordered or quasiperiodic potentials [49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64] and even bosons in interaction-induced disordered or quasiperiodic potentials [65]. The experimental realization of twisted bilayer optical lattices [66] has opened new avenues to investigate interacting bosons in quasicrystals using concrete experimental platforms [67, 68]. More recently, 2D octagonal quasicrystalline optical lattices were directly realized in ultracold-atom experiments by using four light beams in a plane and one beam in the perpendicular direction, and BG with ultracold atoms in the weakly interacting regime was observed [69].
@@
In this work, we study the ground state and dynamical properties of a class of ultracold bosonic systems realized in optical lattices with an aperiodic external potential—a setting directly inspired by recent experiments on twisted bilayer optical lattices [66] and optical quasicrystals [69]. We investigate how the interplay between on-site repulsive interactions and a quasiperiodic potential drives the system between a homogeneous superfluid (SF) phase and a BG phase. At low filling factors, our findings reveal that an increase in the on-site interaction strength induces a delocalization effect, transforming a BG phase—with disconnected SF regions lacking global connectivity—into a robust SF phase characterized by a percolated network of SF clusters. This phase transition is quantitatively characterized by monitoring the percolation probability (see Fig. 1). At higher filling factors, our results uncover a striking reentrant behavior (see Fig. 2): The system initially undergoes a transition from a BG to an SF phase as the on-site interaction strength increases, but a further increase in the interaction leads to a reversal back to the BG phase. This reentrance is attributed to the interaction-driven rearrangement of particles, wherein a once-percolated SF network eventually fragments into isolated SF islands as repulsive interactions become dominant. Moreover, our examination of the system’s quench dynamics reveals distinct transient behaviors (see Fig. 3): While intraphase quenches produce minimal changes in both the percolation probability and inverse participation ratio (IPR) of the particle-density distribution, interphase quenches lead to pronounced dynamical effects. Specifically, a quench from the SF phase to the BG phase results in an abrupt loss of global SF connectivity, whereas a BG-to-SF quench is marked by oscillatory changes in the percolation probability and a gradual decrease in the IPR, ultimately stabilizing the SF phase. These findings unveil the rich physics of ultracold Bose gases in optical lattices with an aperiodic external potential, offering valuable insights that are highly relevant for state-of-the-art experiments employing twisted bilayer or quasicrystalline optical lattice platforms.
@@ II System and Model
Inspired by the recent experimental progress in realizing twisted bilayer optical lattice [66] and optical quasicrystal [69] for ultracold atoms, we consider a single-component ultracold bosonic system, the Hamiltonian of which takes the form of a “single-layer” Bose-Hubbard model with an additional external on-site potential, i.e.,
@@
where $\hat{n}_{i}=\hat{b}_{i}^{\dagger}\hat{b}_{i}$ is the occupation number operator at site $i$, with $\hat{b}_{i}^{\dagger}\penalty 10000\ (\hat{b}_{i})$ being the corresponding bosonic creation (annihilation) operator; $U$ is the strength of the on-site interaction; and $J$ is the hopping amplitude between nearest-neighbor lattice sites (denoted by $\langle\cdots\rangle$). Here, we focus on the case where the underlying optical lattice is a square lattice because it is the most relevant for the current experimental setups [66, 69]. The last term of $\hat{H}$ is the additional on-site external potential with $M_{i}$ being its strength at site $i$. In the experimental setup that realizes the twisted bilayer optical lattice [66], the on-site external potential $M_{i}$ takes into account the influence of the lattice on the other layer in the large-detuning limit, where the bilayer system can be effectively described by the single-layer model [70, 66]. While in the experimental set-up that realizes the optical quasicrystal [69], $M_{i}$ simply corresponds to the additional optical lattice potential felt by the atoms. In this regard, the physics of the model Hamiltonian (1) that will be discussed in the following is relevant for ultracold twisted bilayer optical lattice systems [66] and optical quasicrystal systems [69] in experiments.
@@
where $M_{r}$ is the strength of the external potential, $i_{x}$ ($i_{y}$) is the index of the $i\textrm{th}$ site of the square lattice along the $x$ ($y$) direction, and we choose $\theta=45\text{\textdegree}$. This corresponds to the case in bilayer systems with a twist angle of $45\text{\textdegree}$ [66] or the one in the optical quasicrystal system with the optical lattice formed by superimposing four independent one-dimensional lattices in the $x$-$y$ plane at a $45\text{\textdegree}$ angle [69].
The first two terms of the Hamiltonian (1) assume the form of the conventional Bose-Hubbard model [39]; they favor two homogeneous phases, namely, the homogeneous SF phase (for large $J/U$) and the Mott insulator phase (for small $J/U$ at integer filling factors) that respect the discrete translation symmetry of the underlying lattice. The third term of $\hat{H}$ is an aperiodic or a quasiperiodic external potential that explicitly breaks the lattice translation symmetry. Recent investigations [60] showed that this type of term can give rise to the BG phase that usually arises in disordered bosonic systems [39] due to their aperiodicity. In fact, the interplay between on-site interaction and the quasiperiodic external potential can give rise to much richer physics associated with the BG phase. Particularly, at fixed particle densities that are highly relevant for corresponding experiments, this interplay can give rise to interaction induced reentrant transitions for BG phases and rich quench dynamics as we shall show in the following.
@@ III Results
To study the ground state of the system described by Hamiltonian (1), we used the bosonic Gutzwiller variational method [71, 72, 73]. The variational ground state is assumed to take the site-factorized form $\left|\textrm{GW}\right\rangle=\left|\phi_{1}\right\rangle\otimes\cdots\otimes\left|\phi_{i}\right\rangle\otimes\cdots\otimes\left|\phi_{N_{\textrm{lat}}}\right\rangle$, where each $\left|\phi_{i}\right\rangle$ represents the state at the $i$th lattice site and $N_{\textrm{lat}}$ denotes the total number of sites in the lattice. The individual single-site states $\left|\phi_{i}\right\rangle$ are expanded as a superposition of particle-number eigenstates as $\left|\phi_{i}\right\rangle=\sum_{n=0}^{n_{\textrm{max}}}C_{i.n}\left|n\right\rangle_{i}$, where $n_{\textrm{max}}$ is the maximum number of particles per site and $C_{i.n}$ are variational parameters that are to be determined by minimizing the total energy of the system. In this work, we focus on the ground-state properties of the system at different filling factors $\rho\equiv N/N_{\textrm{lat}}$, with $N$ being the total number of particles in the system. We examine the system at different filling factors to explore how particle density influences the ground-state properties. For the numerical calculations, we set square-lattice size $L=51$, the local maximum truncation number $n_{\textrm{max}}=10$, and we employ open boundary conditions for the numerical results if not otherwise specified in the text.
Figure 1: Phase diagram and representative real-space distributions. (a) Phase diagram for a system of $100$ particles on a $51\times 51$ square lattice, where the black solid line is the phase boundary between BG (top) and SF (bottom) phases. Inset: Dependence of the percolation probability $\mathcal{P}$ (blue curve) and the inverse participation ratio (IPR; red curve) on $U/J$ for points along the dashed line indicated in the phase diagram. The inserted SF order-parameter distributions correspond to the nearby red points marked in the main plot. (b) Real-space distributions of the particle density $\left\langle\hat{n}_{i}\right\rangle$, the amplitude of the SF order parameter $|\langle\hat{b}_{i}\rangle|$, and the discrete field $S_{i}$ near the phase boundary ($M_{r}/J=5.9$). See text for more details.
@@ III.1 Transitions between superfluids and Bose glasses
For relatively low filling factors, the typical properties of the system are summarized in Fig. 1, which shows the phase diagram of the system at $\rho\approx 0.04$, along with the particle-density distribution $\left\langle\hat{n}_{i}\right\rangle$, the SF order parameter distribution $\langle\hat{b}_{i}\rangle$, and the discrete field distribution $S_{i}$. The latter is used to distinguish between SF and BG phases and will be discussed in detail later. As shown in the bottom left inset of Fig. 1(a), in the low-aperiodic-potential regime, as expected from the homogeneous Bose-Hubbard model, the system is in the SF phase, with a nonzero SF order parameter $\langle\hat{b}_{i}\rangle$ that is distributed relatively evenly across the entire lattice. In contrast, in the high-aperiodic-potential regime, as shown in the top inset of Fig. 1(a), the system maintains a nonzero SF order parameter $\langle\hat{b}_{i}\rangle$, but its distribution exhibits a qualitatively different structure. In particular, although the SF order parameter is nonzero in various regions, these regions are “disconnected” from each other, indicating a transition to a new phase: the BG phase, which is absent in conventional homogeneous systems of lattice bosons.
@@
To quantitatively distinguish the SF phase from the BG phase, we employ the percolation method [74, 75, 76, 77, 60], a widely used technique for studying phase transitions in disordered or quasidisordered systems. The key quantity used to differentiate the SF and BG phases is the percolation probability $\mathcal{P}$, with $\mathcal{P}=0$ indicating no connected SF region spanning the entire system (signifying the BG phase) and $\mathcal{P}>0$ indicating the presence of a percolated SF region (signifying a SF phase spanning the entire system). In this work, we adopt the method introduced in [77, 60], where $\mathcal{P}$ is calculated using the corresponding discrete field $S_{i}$ [77], defined as
$S_{i}=\begin{cases}0&\textrm{if}\textrm{ }I-\gamma_{n}\leq\left\langle\hat{n}_{i}\right\rangle\leq I+\gamma_{n},\\ 1&\textrm{otherwise},\end{cases}$ (3)
where $\gamma_{n}=5\times 10^{-3}$ is the threshold used in the numerical calculations and $I=0,1,2,\ldots$ denotes non-negative integer values. This discrete field encapsulates the percolation properties of the SF regions in the system (see Appendix A for technical details on extracting $\mathcal{P}$ from $S_{i}$). For example, in the last row of Fig. 1(b), red sites represent various connected SF regions that do not span the entire system, while blue sites correspond to connected SF regions that do span the entire system, forming a percolated SF region.
@@
From Fig. 1(a), we observe that tuning the on-site interaction strength $U$ can drive the transition between the SF and BG phases in the high-aperiodic-potential regime. As indicated by the horizontal arrow in Fig. 1(a) and the corresponding distributions for $\left\langle\hat{n}_{i}\right\rangle$, $\langle\hat{b}_{i}\rangle$, and $S_{i}$ in Fig. 1(b) at a fixed aperiodic potential strength $M_{r}$, the system transitions from the BG phase with $\mathcal{P}=0$ at relatively small $U$ to the SF phase with a finite percolation probability $\mathcal{P}$ at larger $U$ [see the bottom right inset of Fig. 1(a)]. This transition is driven by a “delocalization” effect of the repulsive on-site interaction: Under low-filling conditions ($\rho<1$), bosons redistribute from highly occupied sites to nearby vacant or weakly occupied sites, broadening the density profile. By contrast, at integer filling, the standard localization effect of strong interactions can also manifest in this system, driving it into a Mott insulator (see Fig. 5 and Appendix B for details). As seen in the bottom right inset of Fig. 1(a), the IPR for the normalized density distribution $\widetilde{\left\langle\hat{n}_{i}\right\rangle}\equiv\left\langle\hat{n}_{i}\right\rangle/\sum_{i=1}^{N_{\textrm{lat}}}\left\langle\hat{n}_{i}\right\rangle$ directly characterizes the extent of localization in the system, i.e., $\textrm{IPR}\equiv\sum_{i=1}^{N_{\textrm{lat}}}\widetilde{\left\langle\hat{n}_{i}\right\rangle}^{2}$. The IPR decreases as $U$ increases, reflecting the delocalization of the particles. At relatively small $U$, the SF regions are localized [see the left most plot for $S_{i}$ in Fig. 1(b)]. As $U$ increases, the SF regions expand, and at large enough $U$, they can merge, forming a percolated SF region that drives the system into the SF phase.
@@
Figure 2: Phase diagram and representative real-space distributions. (a) Phase diagram for a system of $196$ particles on a $51\times 51$ square lattice, where the black solid line is the phase boundary between BG (upper) and SF (lower) phases. Inset: Dependence of the percolation probability $\mathcal{P}$ (blue curve) and IPR (red curve) on $U/J$ for points along the dashed line indicated in the phase diagram. (b) Real-space distributions of the particle density $\left\langle\hat{n}_{i}\right\rangle$, the amplitude of the SF order parameter $|\langle\hat{b}_{i}\rangle|$, and the discrete field $S_{i}$ near the phase boundary ($M_{r}/J=8$). See text for more details.